hi, I am very lost and I don't know how to make a word problem with these equations y=3x-2 and x-y=4 can you please help

Answer by Alan3354(69443)

hi, I am very lost and I don't know how to make a word problem with these equations y=3x-2 and x-y=4 can you please help
-----------
Why do you want to make a "word problem?"
I think the object is to solve the 2 equations for x & y.
------
y=3x-2
x-y=4
------
Since y is solved for in the 1st eqn, sub for y in the 2nd
x - y = 4
x - (3x-2) = 4
x- 3x + 2 = 4
-2x = 2
x = -1
===============
y = 3x-2
y = 3*(-1) - 2
y = -5